# S2R-HDR: A Large-Scale Rendered Dataset for HDR Fusion

### v1

@@ A.6 Additional Dataset Scale Comparison
Table A4: Statistical analysis of data scenarios, time of day, and indoor/outdoor distribution.
@@ B.1 Motion Materials
As demonstrated in Fig. A4, the S2R-HDR dataset comprises three principal categories of motion materials: (a) human subjects with a comprehensive coverage of appearance variations, including garment diversity and gender attributes; (b) vehicular objects incorporating distinct transportation modalities with differential motion patterns; and (c) zoological specimens exhibiting biologically plausible locomotion characteristics. These motion materials are designed for integration into environmental contexts to facilitate dynamic motion synthesis.
Figure A4: Illustration of motion materials.
@@ B.2 High Dynamic Range Environments
As illustrated in Fig. A5, the S2R-HDR dataset presents a collection of high dynamic range environments encompassing both indoor and outdoor configurations. Through systematic utilization of Unreal Engine 5’s Lumen global illumination system, we achieve precise control over environmental lighting parameters. This technical capability enables physics-based synthesis of illumination scenarios spanning three critical lighting regimes: daylight, twilight, and night.
Figure A5: Illustration of high dynamic range environments.
Figure A6: Illustration of image examples of our S2R-HDR.
@@ B.4 HDR Dataset Evaluation Metrics
To quantitatively assess the superiority of our dataset compared to real-world datasets, we employ seven evaluation metrics whose detailed definitions are provided in Tab. A3. Specifically, FHLP and EHL measure the extent of HDR. SI, CF and stdL quantify intra-frame diversity. ALL and DR evaluate overall style.
@@ B.5 Scene and Motion Distributions
Our dataset comprehensively encompasses diverse motion patterns, varied environments, and heterogeneous environmental illumination conditions. The distribution of different categories across the total collection of 24,000 images is detailed in Tab. A4.
@@ B.6 S2R-HDR Image Examples
As shown in Fig. A6, we present additional image examples of S2R-HDR.

### v2

@@ A.12 Test-time Adaptation Runtime
Table 19: Statistical analysis of data scenarios, time of day, and indoor/outdoor distribution.
@@ B.1 Motion Materials
As demonstrated in Figure 14, the S2R-HDR dataset comprises three principal categories of motion materials: (a) human subjects with a comprehensive coverage of appearance variations, including garment diversity and gender attributes; (b) vehicular objects incorporating distinct transportation modalities with differential motion patterns; and (c) zoological specimens exhibiting biologically plausible locomotion characteristics. These motion materials are sourced from two origins: (1) manually created motion sequences by authors, and (2) pre-defined motion patterns from the Unreal Engine 5 materials we purchased, both of which are designed for integration into environmental contexts to facilitate dynamic motion synthesis.
Figure 14: Illustration of motion materials.
@@ B.2 High Dynamic Range Environments
As illustrated in Figure 15, the S2R-HDR dataset presents a collection of high dynamic range environments encompassing both indoor and outdoor configurations. Through systematic utilization of Unreal Engine 5’s Lumen global illumination system, we achieve precise control over environmental lighting parameters. This technical capability enables physics-based synthesis of illumination scenarios spanning three critical lighting regimes: daylight, twilight, and night.
Figure 15: Illustration of high dynamic range environments.
Figure 16: Illustration of image examples of our S2R-HDR.
@@ B.4 HDR Dataset Evaluation Metrics
To quantitatively assess the superiority of our dataset compared to real-world datasets, we employ seven evaluation metrics whose detailed definitions are provided in Table 18. Specifically, FHLP and EHL measure the extent of HDR. SI, CF and stdL quantify intra-frame diversity. ALL and DR evaluate overall style.
@@ B.5 Scene and Motion Distributions
Our dataset comprehensively encompasses diverse motion patterns, varied environments, and heterogeneous environmental illumination conditions. The distribution of different categories across the total collection of 24,000 images is detailed in Table 19.
@@ B.6 S2R-HDR Image Examples
As shown in Figure 16, we present additional image examples of S2R-HDR.
B.7 Further Application of S2R-HDR
Leveraging the advanced rendering capabilities of Unreal Engine 5, we have augmented our customized rendering pipeline with the capability to render multiple specialized image data types. Beyond producing standard output images, the system simultaneously generates four distinct auxiliary data modalities: optical flow fields, depth maps, diffuse reflectance information, and surface normal vectors. In other words, as shown in Figure 17, each frame has its corresponding four additional auxiliary information.
Currently, our dataset is still limited to the HDR fusion task. The provision of such comprehensive supplementary data aims to broaden the utility of S2R-HDR, facilitating its extension to diverse application domains beyond HDR imaging.
Figure 17: Different image types provided in S2R-HDR.
Appendix C The Use of Large Language Models
We only utilize LLMs to polish writing and correct grammar.